A 36 Ω resistor is connected to four 1.5-Vbatteries.

Part A

What is the joule heat loss per minute in the resistor if the batteries are connected in series?

Part B

What is the joule heat loss per minute in the resistor if the batteries are connected in parallel?
Answer #1

Given,

R = 36 Ohm ; V = 1.5 V

A)when the batteries are connected in series

V = 1.5 V

P = V I = V^2/R = 1.5^2/36 = 0.0625 W = 3.75 J/min

Hence, P = 3.75 J/min

B)V = 4 x 1.5 = 6 V

P = 6^2/36 = 1 W

1W = 1 J/s x 60 s/min = 60 J/min

Hence, P = 60 J/min